For Ho: u=90, Ha: u>90, n=150, ybar follows normal distribution, test statistic = Zo= 1.37, a= 0.05, find the critical value, p-value and give conclusion.
Answer #1

solution:

Test statistic = 1.37

P(z >1.37 ) = 1 - P(z < 1.37) = 1 - 0.9147=0.0853 using z table

P-value = 0.0853

=0.05

critical value Z =Z0.05=-1.645